Please help me solve this question: The first three terms of an arithmetic progression are -1049, -1046, -1043.
Which term of this progression will first be positive?

Answer by richwmiller(17219)

an = a + (n - 1)d
0 < -1049 + (n - 1)3
1049<3*(n-1)
350 351 an = a + (n - 1)d
A351 = -1049 + 3(350)
A351=-1049+1050
A351=1
Answer: A351 is the first positive term.